# Ptychoformer: A Physics-Guided Deep Learning Framework for Ptychographic Imaging

### v1

@@ Abstract
Ptychographic imaging confront limitations in applying deep learning (DL) for retrieval from diffraction patterns. Conventional neural architectures are optimized for natural images, overlooking the unique physical characteristics of diffraction data, including radial intensity decay and coherent information distributed in concentric rings. In this paper, we present Ptychoformer, a physics-guided DL framework for ptychographic imaging that aligns attention mechanisms and feature extraction with these diffraction physics properties through introducing a dual-branch architecture which accounts for both local and non-local dependencies from the patterns. It consists of a Polar Coordinate Attention (PCA) mechanism that is inspired by the Ewald construction in X-ray crystallography to enhance high-frequency component fidelity. Experimental results demonstrate Ptychoformer’s superior performance across both simulated and real data in preserving fine details and suppressing artifacts. On simulated data, Ptychoformer achieves up to 5.4% higher PSNR and 4.2% higher SSIM for amplitude retrieval compared to existing methods. For real experimental data, it demonstrates up to 12.5% higher PSNR and 31.3% higher SSIM for amplitude retrieval. Notably, Ptychoformer maintains robust performance under limited training data and low overlap ratios, outperforming existing models.
@@
Ptychography, Physics-Guided Deep Learning, Inverse Imaging Problems, Transformer.
@@ I Introduction
Addressing the inherent limitations of existing methods requires a fundamental reconsideration of deep learning architectures for ptychographic retrieval, with a focus on optimizing neural networks for representing diffraction patterns in reciprocal space. This insight guided the development of Ptychoformer, a novel framework that bridges frequency-domain diffraction patterns and real-space images using physics-informed deep learning techniques. Drawing inspiration from the Ewald construction in X-ray crystallography, as depicted in Fig. 1(b), treating 2D diffraction patterns as projections of the intersections between the Ewald sphere and the reciprocal lattice onto the detector plane, as shown in the left of Fig. 1. This insight revealed that diffraction patterns are inherently defined in spherical rather than planar geometry. Consequently, we developed Polar Coordinate Attention (PCA), an attention mechanism leveraging polar coordinates that naturally aligns with diffraction physics. This polar coordinate-based attention mechanism reframes real-space structural priors in the diffraction context, mapping scattering vector magnitude to $r$ and angular information to $\theta$, thus capturing both radial intensity attenuation and angular coherence. Recognizing the multi-scale and multi-frequency nature of diffraction patterns, as shown in the right side of Fig. 1, Ptychoformer employs a dual-branch architecture combining a Local Dependencies Branch constructed from standard ViT blocks with a NonLocal Coherence block containing our designed PCA. This design balances the capture of long-range and local dependencies in diffraction patterns. Ptychoformer’s architecture intrinsically aligns with ptychography physics, creating a more natural correspondence between model operations and underlying physical processes.
This structural redesign based on physical insights enables our model to more effectively extract and utilize global coherent information and long-range dependencies in diffraction data, thus offering a physics-informed approach to ptychographic retrieval. Our comprehensive experiments on both simulated and real datasets demonstrate Ptychoformer’s significant advantages over existing state-of-the-art methods, particularly in high-frequency artifact suppression and data efficiency. Under challenging conditions including low overlap ratios, and limited training data, our model maintains superior retrieval quality, achieving up to 36.1% reduction in Mean Squared Error (MSE) and 12.03% increase in Structural Similarity Index Measure (SSIM) compared to the best-performing existing end-to-end methods. These improvements could significantly enhance the applicability of ptychography in time-sensitive or radiation-sensitive imaging scenarios across various scientific disciplines.
@@
  • We present Ptychoformer, an end-to-end deep learning framework addressing limitations of existing DL-based ptychography methods. Unlike conventional neural architectures optimized for spatial-domain natural images, our approach incorporates designs aligned with diffraction physics.
  • We propose the PCA mechanism, a physics-inspired approach that decomposes attention into radial and angular components, enabling efficient extraction of coherent information across frequency bands in diffraction patterns.
  • Extensive experiments on both simulated and real datasets show that the proposed method achieves high-fidelity and superior artifact reduction, while maintaining efficiency under challenging conditions including noisy environments and limited data acquisition scenarios.
@@ II-A1 Pre-processing
Integrating deep learning with iterative algorithms improves initial estimates. The physics-informed automatic differentiation ptychography (ADP) framework uses pre-trained autoencoders to map high-dimensional image data to a low-dimensional latent space [15], while the double deep image prior (DDIP) method reduces the optimization parameter space [16], both enhancing convergence rates and noise robustness.
@@ II-A2 Post-processing
Neural networks refine reconstructions from traditional algorithms, including enhancing a single iteration of the Gerchberg-Saxton-Fienup (GSF) algorithm to improve spatial resolution and reduce artifacts [5], showing promise in live cell imaging.
II-A3 End-to-End
End-to-end methods directly map diffraction patterns to complex object functions, bypassing iterative processes. Examples include PtychoNN with a modified U-Net and two-branch decoder for amplitude and phase [9], PtyNet with group convolution and Leaky ReLU for efficiency [11], and DPI using a traditional U-Net with skip connections [10].
Despite improved retrieval speed, these methods face challenges to preserve high-frequency information and the global coherence of diffraction patterns, as CNNs and standard Vision Transformers (ViTs) are tailored for real-space images with local correlations. Recent studies using ViTs for ptychographic retrieval leverage self-attention for long-range dependencies [17], but applying ViT architectures to frequency-domain data is limited by radial intensity decay and concentric information distribution. Addressing these frequency-domain challenges, including processing global coherence and preserving high frequencies, remains crucial for advancing retrieval quality and efficiency, especially with limited data or low overlap ratios.
II-B ViT-based Methods
@@
Figure 2: The proposed Ptychoformer architecture for ptychographic imaging. It features dual branches: a Local Dependencies Branch with standard ViT blocks, and a NonLocal Coherence Branch with PolarCoordinate Attention mechanism. The model processes logarithmically mapped diffraction patterns, combining features from both branches before decoding into amplitude and phase images with physical constraints applied.
Diffraction patterns, governed by wave optics, approximate the Fourier transform of an object’s transmission function in far-field conditions [18], resulting in reciprocal space features like sparse representations and concentric rings [19]. Accurate image retrieval and phase retrieval require effectively capturing both sparsity and long-range dependencies within these patterns. Traditional CNNs, though effective for natural images [20, 21], are limited with the global nature of diffraction phenomena due to their local receptive fields. Vision Transformers (ViTs), introduced by Dosovitskiy et al. [22], offer a promising alternative with self-attention mechanisms that capture global features and long-range dependencies [23], outperforming CNNs in various computer vision tasks [24, 25] and extending to fields like medical image segmentation [24, 26], climate modeling [27], and molecular structure prediction [28]. However, applying ViTs to phase retrieval in CDI is challenging due to the lack of CNN-like inductive biases such as translational invariance and local connectivity, coupled with often insufficient data in practical phase imaging. For example, PtychoDV [17] uses ViTs for feature extraction from diffraction patterns but shows inferior retrieval performance compared to CNN-based U-Net architectures without iterative post-processing. This gap highlights the potential for developing hybrid methods that leverage transformers’ global feature extraction while addressing data limitations in ptychography.
@@ III-A Problem Formulation
where $\psi_{j}(\mathbf{r})=P(\mathbf{r}-\mathbf{r}_{j})\cdot O(\mathbf{r})$ is the exit wave function, $\mathcal{F}$ denotes the Fourier transform, $I_{j}(\mathbf{q})$ is the measured diffraction intensity, and $\phi_{j}(\mathbf{q})$ is the phase of the diffraction pattern.
To account for real-world factors, we model the diffraction intensity as:
$\begin{split}I_{j}(\mathbf{q})=&\mathcal{P}_{\eta}\left(\eta\left|\mathcal{F}\left\{\mu P(\mathbf{r}-\mathbf{r}_{j}-\delta\mathbf{r}_{j})O(\mathbf{r})\right.\right.\right.\\ &\left.\left.\left.+(1-\mu)P(\mathbf{r}-\mathbf{r}_{j}-\delta\mathbf{r}_{j})\mathbb{E}_{\mathbf{r}}[O(\mathbf{r})]\right\}\right|^{2}\right)\\ &+\mathcal{N}(0,\sigma^{2})\end{split}$ (2)
where $\mu$ is the coherence parameter, $\eta$ is the detector efficiency, $\mathcal{P}_{\eta}$ models Poisson shot noise, $\mathcal{N}(0,\sigma^{2})$ represents Gaussian noise, and $\delta\mathbf{r}_{j}$ accounts for positional jitter. The objective function for ptychographic retrieval is defined as:
@@
where $\left\|\mathbf{x}\right\|_{2}^{2}=\sum_{i}x_{i}^{2}$ is the squared Euclidean norm, $\Omega_{1}(O)=\left\|\nabla O\right\|_{1}$ enforces sparsity in the object gradient, $\Omega_{2}(P)=\left\|P-P_{0}\right\|_{2}^{2}$ constrains the probe function to maintain proximity to to an initial estimate $P_{0}$, and $\lambda_{1}$ and $\lambda_{2}$ are regularization weights.
III-B Ptychoformer Architecture
Figure 3: Architectural overview of the PCA mechanism. The diagram illustrates the flow of input features through polar coordinate transformation, query-key-value attention computations, and subsequent operations including polar coordinate weighting and Hadamard product. The process culminates in multi-head attention, followed by concatenation and linear projection to produce the final output.
The Ptychoformer architecture uses a bifurcated structure to map diffraction patterns to a reconstructed complex-valued object function. The input undergoes a logarithmic transformation $I_{\text{log}}(\mathbf{q},j)=\log(1+I_{j}(\mathbf{q}))$ as a preprocessing step to address the large disparity between high and low frequency values in the diffraction patterns.
@@ III-B1 Local Dependencies Branch
The Local Dependencies Branch employs a standard ViT to analyze diffraction patterns, aiming to capture local features and relationships within the pattern. The branch’s structure consists of several key components. Initially, a patch extraction mechanism segments the diffraction pattern into smaller sections. These segments then undergo reshaping to facilitate subsequent processing. Following this, an encoding step is applied to each patch, designed to emphasize relevant features. The encoded patches are then processed through a series of ViT blocks. Each ViT block $\mathcal{B}_{l}$ in $\mathcal{B}_{L}$ is defined as:
$\begin{split}\mathcal{B}_{l}(\mathbf{X})&=\mathbf{X}+\mathcal{FFN}(\mathcal{LN}(\mathbf{X}+\mathcal{MSA}(\mathcal{LN}(\mathbf{X}))))\\ \mathcal{MSA}(\mathbf{X})&=\text{Concat}(\text{head}_{1},\ldots,\text{head}_{h})\mathcal{W}^{O}\\ \text{head}_{i}&=\text{Attention}(\mathbf{X}\mathcal{W}^{Q}_{i},\mathbf{X}\mathcal{W}^{K}_{i},\mathbf{X}\mathcal{W}^{V}_{i})\end{split}$ (4)
where $\mathcal{LN}$ denotes layer normalization, $\mathcal{MSA}$ is multi-head self-attention with $h$ heads, $\mathcal{FFN}$ is a position-wise feed-forward network, and $\mathcal{W}^{Q}_{i},\mathcal{W}^{K}_{i},\mathcal{W}^{V}_{i},\mathcal{W}^{O}$ are learnable weight matrices. Despite apparent disruption, patch extraction facilitates low-frequency information capture, especially from central patches containing crucial low-frequency components.
III-B2 NonLocal Coherence Branch
The NonLocal Coherence Branch operates directly on the full-resolution input, intentionally omitting patch extraction to preserve the global structure of the diffraction pattern. It replaces the multi-head self-attention in standard ViT blocks with PCA, and the radial information introduced by PCA eliminates the need for position embeddings.
While traditional self-attention mechanisms are effective for capturing long-range dependencies, they are not optimized for ptychographic data. Designed for natural images, they incorporate spatial positional encodings $\text{PE}(p)$ that can bias attention towards local regions: $A_{pq}=\text{softmax}(\frac{(x_{p}+\text{PE}(p))^{T}(x_{q}+\text{PE}(q))}{\sqrt{d_{k}}})$. This provides structural priors suited for conventional visual tasks but may not be optimal for frequency-domain diffraction patterns. To address this limitation, we introduce the PCA mechanism. PCA transforms spatial coordinates into polar coordinates, shifting the structural priors from Cartesian spatial locality to the concentric information distribution inherent in frequency-domain diffraction patterns. This transformation enables PCA to better capture the radial and angular dependencies characteristic of ptychographic data. The PCA layer $\mathcal{PCA}:\mathbb{R}^{B\times H\times W\times C}\to\mathbb{R}^{B\times H\times W\times C^{\prime}}$ is defined as:
$\mathcal{PCA}(\mathbf{X})=\text{Reshape}(\mathcal{F}(\text{Flatten}(\mathbf{X})))$ (5)
$\mathcal{F}(\mathbf{X})=\text{Concat}(\text{Head}_{1},\text{Head}_{2},\dots,\text{Head}_{h})\mathcal{W}_{O}$ (6)
$\text{Head}_{i}=\text{softmax}(\mathcal{A}r^{i}(\mathbf{X})+\mathcal{A}\theta^{i}(\mathbf{X}))\mathcal{V}^{i}(\mathbf{X})$ (7)
$\mathcal{A}r^{i}(\mathbf{X})=\frac{\mathcal{Q}^{i}(\mathbf{X})\mathcal{K}^{i}(\mathbf{X})^{T}}{\sqrt{d_{k}}}\odot\mathbf{R}$ (8)
$\mathcal{A}\theta^{i}(\mathbf{X})=\frac{\mathcal{Q}^{i}(\mathbf{X})\mathcal{K}^{i}(\mathbf{X})^{T}}{\sqrt{d_{k}}}\odot\mathbf{\Theta}$ (9)
where $\mathcal{Q}^{i},\mathcal{K}^{i}\in\mathbb{R}^{N\times d_{k}}$, $\mathcal{V}^{i}\in\mathbb{R}^{N\times d_{v}}$ are the linear projections for queries, keys, and values for the $i$-th head, $d_{k}$ is the dimension of the keys, $\mathbf{R},\mathbf{\Theta}\in\mathbb{R}^{N\times N}$ are the normalized radial and angular coordinate matrices, respectively, $\odot$ denotes the Hadamard product (element-wise multiplication), and $\mathcal{W}_{O}\in\mathbb{R}^{hd_{v}\times C^{\prime}}$ is the output projection matrix. Here, $N=H\times W$ is the number of spatial positions. The coordinate matrices are computed as:
$R_{ij}=\frac{\sqrt{(i-W/2)^{2}+(j-H/2)^{2}}}{\sqrt{(W/2)^{2}+(H/2)^{2}}}$ (10)
$\Theta_{ij}=\frac{\arctan 2(j-H/2,i-W/2)}{\pi}$ (11)
where $i\in[1,W]$ and $j\in[1,H]$ are the horizontal and vertical indices of the spatial positions, respectively. This mechanism decomposes attention into radial and angular components, mirroring the symmetry of diffraction patterns. Operating without patch extraction, this branch preserves global structure. Its polar coordinate attention is designed to extract global coherence information, particularly high-frequency components requiring long-range dependency modeling.

### v2

# A Physics-Inspired Deep Learning Framework with Polar Coordinate Attention for Ptychographic Imaging

@@ Abstract
Ptychographic imaging confronts inherent challenges in applying deep learning for phase retrieval from diffraction patterns. Conventional neural architectures, both convolutional neural networks and Transformer-based methods, are optimized for natural images with Euclidean spatial neighborhood-based inductive biases that exhibit geometric mismatch with the concentric coherent patterns characteristic of diffraction data in reciprocal space. In this paper, we present PPN, a physics-inspired deep learning network with Polar Coordinate Attention (PoCA) for ptychographic imaging, that aligns neural inductive biases with diffraction physics through a dual-branch architecture separating local feature extraction from non-local coherence modeling. It consists of a PoCA mechanism that replaces Euclidean spatial priors with physically consistent radial-angular correlations. PPN outperforms existing end-to-end models, with spectral and spatial analysis confirming its greater preservation of high-frequency details. Notably, PPN maintains robust performance compared to iterative methods even at low overlap ratios — well-suited for high-throughput imaging in real-world acquisition scenarios for samples with consistent structural characteristics.
@@
Ptychography, Physics-Inspired Deep Learning, Reciprocal-Space Learning, Transformer.
@@ I Introduction
Addressing the inherent limitations of existing methods requires a fundamental reconsideration of deep learning architectures for ptychographic retrieval, with a focus on optimizing neural networks for representing diffraction patterns in reciprocal space. This insight inspired the development of PPN, a novel framework that bridges frequency-domain diffraction patterns and real-space images using physics-informed deep learning techniques. Drawing inspiration from the Ewald construction in X-ray crystallography, as depicted in Fig. 1(b), treating 2D diffraction patterns as projections of the intersections between the Ewald sphere and the reciprocal lattice onto the detector plane, as shown in the left of Fig. 1. This insight revealed that diffraction patterns are inherently defined in spherical rather than planar geometry. Consequently, we developed Polar Coordinate Attention (PoCA), an attention mechanism leveraging polar coordinates that naturally aligns with diffraction physics. This polar coordinate-based attention mechanism reframes real-space structural priors in the diffraction context, mapping scattering vector magnitude to $r$ and angular information to $\theta$, thus capturing both radial intensity attenuation and angular coherence. Recognizing the multi-scale and multi-frequency nature of diffraction patterns, as shown in the right side of Fig. 1, PPN employs a dual-branch architecture combining a Local Dependencies Branch constructed from standard ViT blocks with a NonLocal Coherence block containing our designed PoCA. This design balances the capture of long-range and local dependencies in diffraction patterns. PPN’s architecture intrinsically aligns with ptychography physics, creating a more natural correspondence between model operations and underlying physical processes.
This structural redesign based on physical insights enables our model to more effectively extract and utilize global coherent information and long-range dependencies in diffraction data, thus offering a physics-informed approach to ptychographic retrieval. Our comprehensive evaluation across simulated and real experimental datasets demonstrates PPN’s multi-faceted advantages: (1) Superior retrieval quality across all metrics against leading end-to-end baselines, particularly in high-frequency preservation as evidenced through power spectral density curves and cross-sectional intensity profiles; (2) Remarkable operational viability - maintaining $<$5% performance degradation at 30% overlap ratio while achieving $>$1,000× faster inference than iterative method when tested on samples without feature distribution shifts, crucial for time-sensitive synchrotron experiments; (3) Unprecedented efficiency with 11× fewer parameters than transformer-based counterparts (6.1M vs 68.9M) and Floating Point Operations Per Second (FLOPs) comparable to lightweight CNNs, enabling deployment in real-world acquisition scenarios. These improvements could significantly enhance the applicability of ptychography in time-sensitive or radiation-sensitive imaging scenarios across various scientific disciplines.
@@
  • We present PPN, a physics-inspired dual-branch framework specifically designed for ptychographic imaging that addresses the geometric mismatch between Euclidean spatial priors and concentric coherent patterns in reciprocal space.
  • We propose the PoCA mechanism that replaces spatial neighborhood priors with radial-angular correlations, achieving superior high-frequency preservation compared to other end-to-end baselines.
  • We demonstrate PPN’s practical advantages with $>$1000× faster inference than iterative method, $<$5% performance degradation at 30% overlap ratio when tested on samples without feature distribution shifts.
@@ II-A1 Pre-processing
Integrating deep learning with iterative algorithms improves initial estimates. The physics-informed automatic differentiation ptychography (ADP) framework uses pre-trained autoencoders to map high-dimensional image data to a low-dimensional latent space [15], while the double deep image prior (DDIP) method reduces the optimization parameter space [16], both enhancing convergence rates and noise robustness. However, these methods still require traditional iterative algorithms.
@@ II-A2 Post-processing
Neural networks refine reconstructions from traditional algorithms, including enhancing a single iteration of the iterative algorithm to improve spatial resolution and reduce artifacts [5] [17]. While these methods significantly improve the quality of reconstructed images, they still rely on initial reconstructions and cannot provide real-time or fully automated solutions.
II-A3 End-to-End with CNNs
End-to-end methods directly map diffraction patterns to complex object functions, bypassing iterative processes. Examples include PtychoNN with a modified U-Net and two-branch decoder for amplitude and phase [9], PtyNet with group convolution and Leaky ReLU for efficiency [11], and DPI using a traditional U-Net with skip connections [10]. Whereas CNNs excel at local feature extraction for natural images[18, 19], their inductive biases prove suboptimal for diffraction patterns requiring global phase coherence. The quadratic phase factors in Fresnel propagation create position-dependent correlations that span the entire detector plane, yet the local receptive fields of CNNs (typically 3×3) cannot span the full radial extent of diffraction rings, causing these physically critical phase relationships between distant pixels to be irrecoverably lost during feature encoding. Furthermore, progressive downsampling in hierarchical architectures systematically discards high-frequency information during feature abstraction. These inherent limitations of CNNs motivate the exploration of attention mechanisms for better long-range dependency modeling.
II-B Transformer-based Methods
@@
Figure 2: The proposed PPN for ptychographic imaging. It features dual branches: a Local Dependencies Branch with standard ViT blocks, and a NonLocal Coherence Branch with Polar Coordinate Attention mechanism. The model processes logarithmically mapped diffraction patterns, combining features from both branches before decoding into individual amplitude and phase reconstructions at each position, which are then separately stitched to generate full field of view images for both amplitude and phase.
Vision Transformers (ViTs) [20] emerged as a promising solution to the limitations of CNNs, demonstrating exceptional capability in capturing long-range dependencies across various domains [21, 22, 23, 24, 25]. Recent adaptations in ptychography, such as PtychoFormer[26] employs a hybrid architecture combining ViT and CNN components, specifically a SegFormer variant adapted for ptychography. While it surpasses traditional CNN methods in reconstruction accuracy, this comes with significant increases in parameter count and training costs. PtychoDV [17] adopts a compromise strategy that utilizes ViTs’ reconstruction results as initial guesses for iterative methods, thereby enhancing the final accuracy of conventional iterative approaches. This indirectly reveals that pure end-to-end deep learning methods still face performance bottlenecks in diffraction reconstruction tasks - the core issue we aim to address in this paper.
While Transformers demonstrate superior capability in capturing global contexts, their direct application to diffraction patterns neglects crucial physical priors inherent in reciprocal space representations. Diffraction patterns, governed by wave optics, approximate the Fourier transform of an object’s transmission function in far-field conditions [27], resulting in reciprocal space features like sparse representations and concentric rings [28]. Accurate image retrieval and phase retrieval require effectively capturing both sparsity and long-range dependencies within these patterns. Although Vision Transformers achieve global modeling through multi-head self-attention mechanisms (MHSA), their Euclidean spatial neighborhood-based attention weight calculation exhibits geometric mismatch with the concentric coherent patterns in reciprocal space characteristic of diffraction patterns. Our focus is to resolve the critical bottleneck of effective information extraction from diffraction patterns and improved reconstruction of high-frequency spatial information, rather than surpassing or replacing the ultimate resolution achieved by iterative algorithms.
@@ III-A Problem Formulation
where $\psi_{j}(\mathbf{r})=P(\mathbf{r}-\mathbf{r}_{j})\cdot O(\mathbf{r})$ is the exit wave function, $\mathcal{F}$ denotes the Fourier transform, $I_{j}(\mathbf{q})$ is the measured diffraction intensity, and $\phi_{j}(\mathbf{q})$ is the phase of the diffraction pattern. To account for real-world factors, we model the diffraction intensity as:
$\begin{split}I_{j}(\mathbf{q})=&\text{Poisson}\Big{(}\eta\Big{|}\mathcal{F}\Big{\{}\mu P(\mathbf{r}-\mathbf{r}_{j}-\delta\mathbf{r}_{j})O(\mathbf{r})\\ &+(1-\mu)P(\mathbf{r}-\mathbf{r}_{j}-\delta\mathbf{r}_{j})\cdot\mathbb{E}_{\mathbf{r}}[O(\mathbf{r})]\Big{\}}\Big{|}^{2}\Big{)}\\ &+\mathcal{N}(0,\sigma^{2})\end{split}$ (2)
where $\mu$ is the coherence parameter, $\eta\in[0,1]$ is the detector efficiency representing the quantum efficiency (probability of detecting each photoelectron), $\text{Poisson}(\lambda)$ denotes the Poisson distribution with rate parameter $\lambda=\eta\cdot\left|\mathcal{F}\left\{\dots\right\}\right|^{2}$, where |{}$\left|\mathcal{F}\left\{\dots\right\}\right|^{2}$ is the theoretical diffraction intensity (in photons/pixel) and their product forms the expected value of the Poisson process, and $\delta\mathbf{r}_{j}$ accounts for positional jitter.
Based on the physical model of the actual imaging process in Eq. (2), we formulate the objective function for ptychographic retrieval in Eq. (3) to effectively recover the object and probe functions:
@@
where $\left\|\mathbf{x}\right\|_{2}^{2}=\sum_{i}x_{i}^{2}$ is the squared Euclidean norm, $\Omega_{1}(O)=\left\|\nabla O\right\|_{1}$ enforces sparsity in the object gradient, $\Omega_{2}(P)=\|P-P_{0}\|_{F}^{2}$ constrains the probe function to maintain proximity to to an initial estimate $P_{0}$, and $\lambda_{1}$ and $\lambda_{2}$ are regularization weights.
III-B PPN Architecture
The proposed architecture employs a bifurcated structure to map diffraction patterns to a reconstructed complex-valued object function. The input undergoes a logarithmic transformation $I_{\text{log}}(\mathbf{q},j)=\log(1+I_{j}(\mathbf{q}))$ as a preprocessing step to address the large disparity between high and low frequency values in the diffraction patterns.
@@ III-B1 Local Dependencies Branch
The Local Dependencies Branch utilizes a standard Vision Transformer (ViT) to analyze diffraction patterns, capturing local features and structural relationships within each pattern. Operating on input tensors $\mathbf{X}\in\mathbb{R}^{B\times H\times W\times C}$ (where $B$ is batch size, $H\times W$ is spatial resolution, and $C$ is channel depth), this branch processes data through several key components:
Step 1. Patch Extraction divides the input image into non-overlapping patches of size $P\times P$, resulting in $\mathbf{X}_{\text{patches}}\in\mathbb{R}^{B\times N_{p}\times(P^{2}\cdot C)}$ where $N_{p}=\frac{H}{P}\times\frac{W}{P}$ represents the total number of patches.
Step 2. Linear Projection maps each patch to a $D$-dimensional embedding space using $\mathbf{X}_{\text{embed}}=\mathbf{X}_{\text{patches}}\mathbf{W}_{E}\in\mathbb{R}^{B\times N_{p}\times D}$ where $\mathbf{W}_{E}\in\mathbb{R}^{(P^{2}C)\times D}$ is the projection matrix.
Step 3. Positional Encoding adds spatial information via $\mathbf{X}_{\text{pos}}=\mathbf{X}_{\text{embed}}+\mathbf{E}_{\text{pos}}$, where the positional encoding matrix $\mathbf{E}_{\text{pos}}\in\mathbb{R}^{N_{p}\times D}$ is broadcasted across the batch dimension to match $\mathbf{X}_{\text{embed}}\in\mathbb{R}^{B\times N_{p}\times D}$.
Step 4. Transformer Blocks with Pre-LN structure process the sequence through $L$ layers according to $\mathbf{X}^{\prime(l)}=\mathcal{MSA}(\mathcal{LN}(\mathbf{X}^{(l-1)}))+\mathbf{X}^{(l-1)}$ and $\mathbf{X}^{(l)}=\mathcal{FFN}(\mathcal{LN}(\mathbf{X}^{\prime(l)}))+\mathbf{X}^{\prime(l)}$. The Multi-Head Self-Attention (MSA) mechanism computes $\mathcal{MSA}(\mathbf{X})=\text{Concat}(\text{head}_{1},...,\text{head}_{h})\mathbf{W}^{O}$ where each attention head captures different relationship patterns in the data, with projection matrices $\mathbf{W}^{Q}_{i},\mathbf{W}^{K}_{i}\in\mathbb{R}^{D\times d_{k}}$ and $\mathbf{W}^{V}_{i}\in\mathbb{R}^{D\times d_{v}}$ (typically $d_{k}=d_{v}=D/h$).
Step 5. Spatial Restoration reshapes the output to $\mathcal{H}_{\text{Local}}\in\mathbb{R}^{B\times(H/P)\times(W/P)\times D}$.
[…]
III-B2 Non-Local Coherence Branch
The Non-Local Coherence Branch employs an architectural framework similar to the Local Dependencies Branch, sharing identical transformer block components while introducing our novel Polar Coordinate Attention (PoCA) mechanism. This specialized attention mechanism processes input at pixel-level granularity ($H\times W$ pixels) without patch extraction, enabling the capture of long-range dependencies across diffraction patterns. The PoCA mechanism systematically encodes physical constraints inherent to diffraction physics, including the $I\propto r^{-4}$ intensity distribution through radial decay weighting, preserves crystallographic symmetry via angular continuity, and adapts to illumination shifts with dynamically learned centering parameters $\alpha_{x}$ and $\alpha_{y}$.
After logarithmic transformation, the single-channel diffraction pattern $\mathbb{R}^{B\times H\times W\times 1}$ is projected to a multi-channel tensor $\mathbb{R}^{B\times H\times W\times C}$ via a learnable $1\times 1$ convolutional layer, where $C$ denotes the embedding dimension. For an input tensor $\mathbf{X}\in\mathbb{R}^{B\times H\times W\times C}$ (where $B$ represents batch size, $H\times W$ spatial resolution), this branch processes data through several key components:
Step 1. Tensor Flattening transforms the input into a sequence representation via $\mathbf{X}_{f}=\text{Reshape}(\mathbf{X},(B,H\times W,C))\in\mathbb{R}^{B\times N\times C}$, where $N=H\times W$ represents the total number of pixels.
Step 2. Query-Key-Value Projection maps the flattened tensor into three separate embedding spaces using:
$\mathcal{Q}^{i}$ $=\mathbf{X}_{f}\mathbf{W}_{Q}^{i}\quad$ $\mathbf{W}_{Q}^{i}$ $\in\mathbb{R}^{C\times d_{k}}$ (4)
$\mathcal{K}^{i}$ $=\mathbf{X}_{f}\mathbf{W}_{K}^{i}\quad$ $\mathbf{W}_{K}^{i}$ $\in\mathbb{R}^{C\times d_{k}}$
$\mathcal{V}^{i}$ $=\mathbf{X}_{f}\mathbf{W}_{V}^{i}\quad$ $\mathbf{W}_{V}^{i}$ $\in\mathbb{R}^{C\times d_{v}}$
where projection matrices $\mathbf{W}_{Q}^{i},\mathbf{W}_{K}^{i}\in\mathbb{R}^{C\times d_{k}}$ and $\mathbf{W}_{V}^{i}\in\mathbb{R}^{C\times d_{v}}$ define the embedding transformations for attention head $i$. The resulting tensors are $\mathcal{Q}^{i},\mathcal{K}^{i}\in\mathbb{R}^{B\times N\times d_{k}}$ and $\mathcal{V}^{i}\in\mathbb{R}^{B\times N\times d_{v}}$, where the matrix multiplication is performed along the feature dimension $C$. Polar coordinate parameterization defines the physical diffraction geometry using:
$c_{x}=\frac{W}{2}+\alpha_{x}\frac{W}{2},\quad c_{y}=\frac{H}{2}+\alpha_{y}\frac{H}{2}$ (5)
$r_{m}=\frac{\log(1+\|\mathbf{p}_{m}-\mathbf{c}\|)}{\log(1+r_{\max})},\quad\theta_{m}=\arctan 2\left(y_{m}-c_{y},x_{m}-c_{x}\right)$ (6)
$\mathbf{\Phi}_{r}^{mn}=\frac{1}{1+|r_{m}-r_{n}|},\quad\mathbf{\Phi}_{\theta}^{mn}=\cos(\theta_{m}-\theta_{n})$ (7)
where $r_{\max}=\sqrt{(W/2)^{2}+(H/2)^{2}}$ ensures normalized radial coordinates, $\mathbf{p}_{m}=(x_{m},y_{m})$ denotes original pixel coordinates, $\mathbf{c}=(c_{x},c_{y})$ represents the learned diffraction center, and $\alpha_{x},\alpha_{y}\in[-0.5,0.5]$ are learnable parameters initialized at 0. The implementation employs logarithmic scaling to compress the dynamic range of radial distances, with angle normalization $\theta_{m}\in[0,2\pi)$ ensuring continuity.
Step 3. Polar Coordinate Modulation incorporates physical constraints by modulating the standard dot-product attention with radial and angular weighting matrices:
$\begin{split}\mathcal{A}_{\text{polar}}^{i}&=\underbrace{\frac{\mathcal{Q}^{i}(\mathcal{K}^{i})^{\top}}{\sqrt{d_{k}}}}_{\text{base attention}}\odot\underbrace{\mathbf{\Phi}_{r}}_{\text{radial decay}}\odot\underbrace{\mathbf{\Phi}_{\theta}}_{\text{angular continuity}}\end{split}$ (8)
where $\mathbf{\Phi}_{r}$, $\mathbf{\Phi}_{\theta}\in\mathbb{R}^{1\times N\times N}\quad\text{(broadcastable to }\mathbb{R}^{B\times N\times N}\text{)}$, $\odot$ represents element-wise multiplication with broadcasting, imposing both radial decay and angular continuity constraints.
Step 4. Attention Application computes each attention head through $\text{Head}_{i}=\text{softmax}\left(\mathcal{A}_{\text{polar}}^{i}\right)\mathcal{V}^{i}$, where the softmax normalization ensures proper probability distribution across the attention weights, and $\mathcal{A}_{\text{polar}}^{i}\in\mathbb{R}^{B\times N\times N}$ is applied to $\mathcal{V}^{i}\in\mathbb{R}^{B\times N\times d_{v}}$.
Step 5. Multi-head Integration combines information from all h attention heads via $\mathcal{G}(\mathbf{X}_{f})=\text{Concat}(\text{Head}_{1},...,\text{Head}_{h})\mathbf{W}^{O}$, with $\mathbf{W}^{O}\in\mathbb{R}^{h\cdot d_{v}\times C^{\prime}}$ serving as the output projection matrix and $\text{Head}_{i}\in\mathbb{R}^{B\times N\times d_{v}}$.
Step 6. Spatial Restoration reshapes the processed output to match the original spatial dimensions through $\mathcal{P}\text{o}CA{}(\mathbf{X})=\text{Reshape}\left(\mathcal{G}(\mathbf{X}_{f}),(B,H,W,C^{\prime})\right)\in\mathbb{R}^{B\times H\times W\times C^{\prime}}$.
[…]
By decomposing attention into radial and angular components, PoCA effectively mimics the Ewald sphere’s intersection with the reciprocal lattice (Fig. 1(b)), where $r_{mn}$ correlates with the reciprocal space resolution of scattering vectors, and $\theta_{mn}$ encodes Bragg angle relationships through angular coherence constraints.